# CuLEmo: Cultural Lenses on Emotion - Benchmarking LLMs for Cross-Cultural Emotion Understanding

### v1

@@ Abstract
NLP research has increasingly focused on subjective tasks such as emotion analysis. However, existing emotion benchmarks suffer from two major shortcomings: (1) they largely rely on keyword-based emotion recognition, overlooking crucial cultural dimensions required for deeper emotion understanding, and (2) many are created by translating English-annotated data into other languages, leading to potentially unreliable evaluation. To address these issues, we introduce Cultural Lenses on Emotion (CuLEmo), the first benchmark designed to evaluate culture-aware emotion prediction across six languages: Amharic, Arabic, English, German, Hindi, and Spanish. CuLEmo comprises 400 crafted questions per language, each requiring nuanced cultural reasoning and understanding. We use this benchmark to evaluate several state-of-the-art LLMs on culture-aware emotion prediction and sentiment analysis tasks. Our findings reveal that (1) emotion conceptualizations vary significantly across languages and cultures, (2) LLMs performance likewise varies by language and cultural context, and (3) prompting in English with explicit country context often outperforms in-language prompts for culture-aware emotion and sentiment understanding. We hope this benchmark guides future research toward developing more culturally aligned NLP systems.
@@
Tadesse Destaw Belay1,2, Ahmed Haj Ahmed3, Alvin Grissom II3, Iqra Ameer4, Grigori Sidorov1, Olga Kolesnikova1, Seid Muhie Yimam5 1Instituto Politécnico Nacional, 3Haverford College, 3Wollo University, 4Pennsylvania State University, 5University of Hamburg, Contact: tadesseit@gmail.com
@@ 1 Introduction
Figure 1: CuLEmo creation pipeline and evaluations of LLMs in emotion and sentiment tasks.
LLMs and agent systems are being employed to interact extensively with humans across various applications such as customer service, healthcare, and education Wang et al. (2024). To facilitate effective interaction, incorporating aspects of cognitive and emotional-social intelligence—the ability to recognize and interpret human emotions Mathur et al. (2024)—can facilitate better interactions with more people. Factors such as age, cultural background, and personal experiences influence how individuals perceive and process information, particularly within subjective NLP tasks. Among these, emotion recognition (ER) and sentiment analysis (SA) are particularly sensitive to language- and culture-specific nuances Plaza-del-Arco et al. (2024).
Natural language frequently encodes emotional information Jim et al. (2024). For example, consider tipping customs in restaurants: in some cultures (e.g. North America), tipping is widely practiced, whereas in China, it is rare, and in Japan, it may even be considered offensive Givi and Galak (2017). Such differences underscore the importance of culturally-aware language technologies.
@@
  1. 1.
    Do LLMs provide culturally aware-emotional responses?
  2. 2.
    Which cultures are more effectively represented in LLMs?
  3. 3.
    Can LLMs identify a country’s culture based on the text describing an event on the prompt used?
  4. 4.
    Does the language of the prompt affect an LLM’s ability for culture-aware emotion understanding?
@@ 2 Related Work
We now review related work in culturally-aware NLP, culture-oriented benchmarks, and cross-lingual study of linguistic emotional expression. Although culture is a complex concept, most definitions of culture encompass people, groups of people, and interactions between individuals and groups Liu et al. (2024c). In NLP, understanding culture is important for the safety and fairness of Large Language Models (LLMs).
@@ 2.2 Emotion Across Languages
Several studies have examined multilingual emotion datasets. Some findings suggest that emotion categories can be preserved through machine translation, for example, by exploring how English-annotated emotion data translate into Finnish, French, German, Hindi, and Italian Kajava et al. (2020); Tahir et al. (2023); Bianchi et al. (2022). This work suggests that the changes in emotion labels often stem from the inherent difficulty of annotation rather than from linguistic differences.
@@ 2.3 Emotions and Cultures
Recent work examines how cultural contexts shape emotional expression across languages. Havaldar et al. (2023a) analyzed embeddings of 271 emotion keywords in English, Spanish, Chinese, and Japanese by projecting into a Valence-Arousal plane using XLM-RoBERTa Reimers and Gurevych (2020) embeddings, revealing that multilingual models embed non-English emotion words differently. Havaldar et al. (2023b) analyzes Pride/Shame as a known cultural difference by prompting GPT-3.5 and GPT-4 to explore how these models handle pride and shame in the USA vs. Japan. They found that GPT-3.5 displayed limited knowledge of culturally specific norms. Ahmad et al. (2024) expands the 19 cultural questions of Havaldar et al. (2023b) to 37 questions and evaluates ChatGPT for the low-resource Hausa language for sentiment analysis. However, these evaluations are limited to 19/37 culturally relevant questions (not diverse and representative data), limited classes (Pride/Shame or positive/negative/mixed), and limited cultures (USA vs. Japan or Hausa).
@@ 3.1 Collecting Cultural Events
We manually crafted scenarios, searched online, and prompted LLMs to gather traditions, events, norms, and actions that elicit culturally different emotions across six target countries (USA, UAE, Germany, Ethiopia, India, and Mexico).We drew inspiration from the work of Havaldar et al. (2023b) to enhance topic diversity. Language representatives are asked to propose events distinct from those of other countries in the form of emotion-oriented questions. Importantly, these events do not contain explicit emotional keywords (as typically seen in traditional emotion datasets). We also referred to the International Survey on Emotion Antecedents and Reactions (ISEAR) (Scherer and Wallbott, 1994) data format, "When I … situations that cause a specific emotion"—a well-known English dataset for emotion analysis consisting of self-reported events from around 3,000 respondents across 37 countries and five continents. Table 1 displays the ten broad categories and the number of questions in each category.
@@
Table 1: Ten broad categories and the number of events (Qn - questions) in the CuLEmo dataset.
@@ 3.2 Human-Adapted Translation
After collecting the events in English, we translated them into five other target languages—Arabic, Amharic, German, Hindi, and Spanish—using Google Translate, followed by native-speaker approvals. Because the questions are simple “How do you feel when … ?” and lack explicit emotion keywords, translation quality did not affect their cultural content. While we acknowledge that other existing works often depend on translating emotion-annotated data from English into other languages with its labels Kajava et al. (2020); Tahir et al. (2023); Bianchi et al. (2022), the translation process we employed in our study is done before any annotation. To ensure correctness, native speakers evaluated the translations. While most translations were acceptable, a few adjustments were made—for example, for Amharic to fix gender references and timing expressions.
@@ 3.3 Language and Cultures Covered
Several factors guided our choice of languages: (1) typological variety (five distinct languages with four different scripts), (2) geographical diversity (eastern vs. western contexts), (3) resource availability (low- vs. high-resource languages), and (4) the availability of native speakers for translation reviews. These considerations helped ensure a broad range of cultural norms and concepts.
@@ 3.4 CuLEmo Annotation
We used Amazon Mechanical Turk (MTurk) for most annotations, ensuring at least five native-speaker annotators per instance from the respective targeted country. We used a customized POTATO Pei et al. (2022) annotation tool for languages lacking sufficient MTurk annotators (e.g., Amharic) and recruited local native speakers who met our criteria. Annotators were fairly compensated \$12/hr. (Prolific [^1] https://www.prolific.com/’s minimum annotation wage is \$9/hr). In cases where no majority vote emerged among five annotators, we collected two additional annotations and used majority voting to finalize the label.
Figure 2: Emotion label distribution across countries/languages.
@@
Pairwise Label Agreements Across Countries. We further examined label differences across countries by calculating pairwise agreements after majority voting (Figure 3). Ethiopia (ET) and the United Arab Emirates (AE) exhibit the highest agreement at 55%, along with Germany (DE) and the United Arab Emirates (AE), both exhibit a high neutral class; see Figure 2. In contrast, Germany (DE) and India (IN) show the lowest agreement, at 29.0%. Labels from India thus tend to diverge from those of other countries.
Figure 3: Pairwise emotion label agreements across countries (ordered by their average agreement). Abbreviations: US = USA, MX = Mexico, DE = Germany, AE = UAE, ET = Ethiopia, and IN = India.
@@ 4.1 Task Formulation
To investigate the emotional comprehension capabilities of LLMs, we examine how they associate different countries’ cultures with their corresponding languages. Specifically, we explore culture-aware emotion understanding via two main tasks: emotion prediction and sentiment analysis. All tested models are instruction-fine-tuned, except for the Aya-expanse model. We also experiment with prompts that do and do not include explicit country context, using the phrase "You live in <<country name>>," (where <<country name>> is one of the six targeted countries: USA, UAE, Ethiopia, Germany, India, and Mexico). Each task is framed as a text generation problem, and the models are evaluated in a zero-shot setting.
@@ 4.2 Model Selection
We evaluate a variety of recent LLMs known for strong performance on standard benchmarks. We aim to include both smaller and medium-sized models, as well as open-source and proprietary options. The models evaluated include:
@@
    Open-Source Models: LLAMA-3 (3.2-3B, 3.1-8B) Dubey et al. (2024); Meta AI (2024), GEMMA (2B, 9B) Team Gemma et al. (2024), Aya (expanse-8b, 101-13B) Üstün et al. (2024); Dang et al. (2024), Ministral (3B, 8B) Mistral AI (2024)
@@
    Proprietary Models: GPT (3.5, 4) OpenAI (2024), GEMINI-1.5 Google Deepmind (2024), Claude (3.5-sonnet, 3-opus) Anthropic AI (2024)
@@ 4.3 Multilingual Prompt Construction
To examine the impact of prompt language on model performance—and to assess each model’s cultural awareness—we design both English and in-language prompts. The instruction, input text, and expected answer are in English when using English prompts. For in-language prompts, all elements are in one of the five target languages—Arabic (AR), German (DE), Amharic (AM), Hindi (HI), or Spanish (ES). Complete examples of our multilingual prompts for both emotion prediction and sentiment analysis are provided in Appendix C. We extract each model’s answer from its generated text using PEDANTS Li et al. (2024).
@@ 5 Result and Analysis
5.1 Culture-Aware Emotion Prediction
@@
Table 2: LLMs’ accuracy for the emotion prediction task. Columns labeled EN/AR/DE/AM/HI/ES show the prompt language for each corresponding culture (e.g., UAE is tested with English and Arabic). The highest-scoring model across English and in-language prompts is highlighted in bold, and the best model for the in-language prompt is underline. * is not an instruction fine-tuned model.
5.2 Culture Representation in LLMs
LLMs Country
US AE DE ET IN MX
Llama-3.2-3B 0.44 0.41 0.36 0.35 0.31 0.46
Llama-3.1-8B 0.62 0.54 0.50 0.51 0.36 0.60
Gemma-2-2B 0.61 0.59 0.68 0.49 0.32 0.57
Gemma-2-9B 0.59 0.54 0.54 0.50 0.41 0.57
Aya-101-13B 0.61 0.57 0.60 0.55 0.38 0.58
Ministral-8B 0.66 0.60 0.56 0.50 0.38 0.59
Claude-3.5-s. 0.56 0.51 0.39 0.49 0.36 0.60
Claude-3-op 0.52 0.40 0.37 0.49 0.36 0.59
Gemini1.5 0.53 0.51 0.39 0.48 0.39 0.59
GPT-4 0.58 0.55 0.46 0.48 0.37 0.63
Table 3: LLMs’ accuracy using English prompts without specifying the country context, "You live in «country name»". The columns list each target country. The bold columns indicate the top three most accurately represented cultures.
Results: According to Table 3, the USA, Mexico, and Germany consistently achieve higher accuracy scores, while the UAE, Ethiopia, and India remain less accurately represented. This suggests that certain cultures may be more prevalent in the underlying training data.
5.3 Does Language Represent Country?
Can LLMs identify one country’s culture based solely on the prompt language? In this experiment, we remove explicit country context (e.g., "You live in <<country name>>") and test if LLMs can infer cultural cues only from the language used in the prompt.
LLMs Language
EN AR DE AM HI ES
Llama-3.2-3B 0.44 0.13 0.19 0.30 0.25 0.37
Llama-3.1-8B 0.62 0.46 0.31 0.18 0.32 0.56
Gemma-2-2B 0.61 0.47 0.56 0.14 0.27 0.52
Gemma-2-9B 0.59 0.59 0.59 0.28 0.34 0.61
Aya-101-13B 0.61 0.34 0.39 0.31 0.30 0.49
Ministral-8B 0.66 0.07 0.67 0.20 0.23 0.34
Claude-3.5-s. 0.56 0.51 0.35 0.53 0.38 0.61
Claude-3-op 0.52 0.45 0.30 0.47 0.37 0.58
Gemini1.5 0.53 0.54 0.42 0.52 0.37 0.65
GPT-4 0.58 0.46 0.45 0.27 0.41 0.64
Table 4: LLMs’ accuracy for in-language prompts without country context. For each language, bold marks the best performance.
Results: Table 4 shows that accuracy drop significantly when country context is omitted. Comparing these scores with Table 2 (where country context is included), we see consistent performance boosts (e.g. +1% in Spanish with GPT-4, +5% in Arabic with Gemini1.5, +6% in German with Ministral, +9% in Amharic with Claude-3.5-sonnet, and +21% in Hindi with GPT-4) when the prompt explicitly names the country. This indicates that language alone does not reliably convey cultural context. Notably, models like Claude-3.5-sonnet and GPT-4 show improvements in Indian and Ethiopian data when country context is specified, while English prompts (without “USA” context) are less affected. Overall, providing country name information remains crucial for accurate culture-aware emotion understanding, especially for less-resourced languages.
5.4 Culture-Aware Sentiment Analysis
For the sentiment analysis experiment, we follow De Bruyne et al. (2019); Mostafazadeh Davani et al. (2022) by grouping emotions into positive (joy), negative (fear, anger, guilt, sadness), and neutral sentiments. Table 5 shows the accuracy of each LLM under this three-class setup.
LLMs USA UAE Germany Ethiopia India Mexico
EN EN AR EN DE EN AM EN HI EN ES
Llama-3.2-3B 0.57 0.41 0.15 0.20 0.24 0.48 0.51 0.60 0.56 0.55 0.60
Llama-3.1-8B 0.65 0.58 0.50 0.49 0.43 0.59 0.43 0.57 0.55 0.64 0.68
Gemma-2-2B 0.68 0.65 0.58 0.66 0.59 0.56 0.46 0.50 0.59 0.66 0.66
Gemma-2-9B 0.55 0.57 0.63 0.49 0.56 0.62 0.50 0.57 0.61 0.63 0.68
Aya-expanse-8b* 0.45 0.33 0.26 0.31 0.27 0.39 0.50 0.45 0.58 0.45 0.57
Aya-101-13B 0.65 0.64 0.55 0.69 0.45 0.55 0.53 0.52 0.56 0.61 0.63
Ministral-8B 0.70 0.66 0.06 0.60 0.73 0.59 0.46 0.54 0.24 0.65 0.53
Claude-3-opus 0.63 0.60 0.56 0.46 0.35 0.67 0.58 0.59 0.61 0.72 0.75
Gemini1.5-flash 0.63 0.63 0.63 0.48 0.50 0.63 0.60 0.60 0.53 0.69 0.71
GPT-4 0.67 0.61 0.54 0.48 0.51 0.62 0.49 0.58 0.62 0.72 0.75
Table 5: LLMs’ accuracy for culture-aware sentiment analysis (positive/negative/neutral). Columns correspond to English (EN) prompts and the native language (e.g., Arabic for UAE). The highest-scoring model across English and in-language prompts is highlighted in bold, and the best model for the in-language prompt is underlined. * is not an instruction fine-tuned model.
Results: Table 5 shows better overall performance on sentiment analysis compared to fine-grained emotion classification. For example, GPT-4 gains notable accuracy (e.g., +22% for Hindi). Similarly, the highest score appears for Mexican culture (Spanish) with Claude-3-opus and GPT-4, each at 75%. By contrast, the non-instruction-tuned Aya-expanse-8b struggles more as it is not instruction fine-tuned. Smaller models like Gemma-2-2B remain competitive with proprietary systems. Nonetheless, performance drops persist for Amharic and Hindi, reflecting the continued challenges of culture-aware tasks in lower-resource contexts.
Questions US AE DE ET IN MX
@@ 5.4 Culture-Aware Sentiment Analysis
Table 6: Examples of predictions from GPT-4 model. The first row under the country-code columns in each section represents the gold-label emotions, while the second row displays the predicted label. The predictions are using country context, by adding explicitly You live <<country name>> in the prompt

### v2

@@ Abstract
NLP research has increasingly focused on subjective tasks such as emotion analysis. However, existing emotion benchmarks suffer from two major shortcomings: (1) they largely rely on keyword-based emotion recognition, overlooking crucial cultural dimensions required for deeper emotion understanding, and (2) many are created by translating English-annotated data into other languages, leading to potentially unreliable evaluation. To address these issues, we introduce Cultural Lenses on Emotion (CuLEmo), the first benchmark designed to evaluate culture-aware emotion prediction across six languages: Amharic, Arabic, English, German, Hindi, and Spanish. CuLEmo comprises 400 crafted questions per language, each requiring nuanced cultural reasoning and understanding. We use this benchmark to evaluate several state-of-the-art LLMs on culture-aware emotion prediction and sentiment analysis tasks. Our findings reveal that (1) emotion conceptualizations vary significantly across languages and cultures, (2) LLMs performance likewise varies by language and cultural context, and (3) prompting in English with explicit country context often outperforms in-language prompts for culture-aware emotion and sentiment understanding. The dataset [^1] https://huggingface.co/llm-for-emotion and evaluation code [^2] https://github.com/llm-for-emotion/culemo is available.
@@
Tadesse Destaw Belay1,3, Ahmed Haj Ahmed2, Alvin Grissom II2, Iqra Ameer4, Grigori Sidorov1, Olga Kolesnikova1, Seid Muhie Yimam5 1Instituto Politécnico Nacional, 2Haverford College, 3Wollo University, 4Pennsylvania State University, 5University of Hamburg, Contact: tadesseit@gmail.com
@@ 1 Introduction
Figure 1: CuLEmo dataset creation pipeline and evaluations of LLMs in emotion and sentiment tasks.
LLMs and agent systems are employed to interact extensively with humans across applications such as customer service, healthcare, and education Wang et al. (2024). To facilitate effective interaction, incorporating aspects of cognitive and emotional-social intelligence—the ability to recognize and interpret human emotions Mathur et al. (2024)—can facilitate better interactions with more people. Factors such as age, cultural background, and personal experiences influence how individuals perceive and process information, particularly within subjective NLP tasks. Among these, emotion recognition (ER) and sentiment analysis (SA) are particularly sensitive to language- and culture-specific nuances Plaza-del-Arco et al. (2024).
Natural language frequently encodes emotional information Jim et al. (2024). For example, consider tipping customs in restaurants: in some cultures (e.g., North America), tipping is widely practiced, whereas in China, it is rare, and in Japan, it may even be considered offensive Givi and Galak (2017). Such differences underscore the importance of culturally-aware language technologies.
@@
  • RQ1. Do LLMs provide culturally aware-emotional responses?
  • RQ2. Which cultures are more effectively represented in LLMs?
  • RQ3. Can LLMs identify a country’s culture based on the text describing an event in the prompt?
  • RQ3. Does the language of the prompt affect the ability of LLMs for culture-aware emotion understanding?
@@ 2 Related Work
We now review related work in culturally-aware NLP, culture-oriented benchmarks, and cross-lingual study of linguistic emotional expression. Although culture is a complex concept, most definitions of culture encompass people, groups of people, and interactions between individuals and groups Liu et al. (2024c). Understanding culture is important for the safety and fairness of LLMs.
@@ 2.2 Emotion Across Languages
Several studies have examined multilingual emotion datasets. Some findings suggest that emotion categories can be preserved through machine translation, for example, by exploring how English-annotated emotion data translate into Finnish, French, German, Hindi, and Italian Kajava et al. (2020); Tahir et al. (2023); Bianchi et al. (2022). These works suggest that the changes in emotion labels often stem from the inherent difficulty of annotation rather than from linguistic differences.
@@ 2.3 Emotions and Cultures
Recent work examines how cultural contexts shape emotional expression across languages. Havaldar et al. (2023a) analyze embeddings of 271 emotion keywords in English, Spanish, Chinese, and Japanese by projecting into a Valence-Arousal plane using XLM-RoBERTa Reimers and Gurevych (2020) embeddings, finding that multilingual models embed non-English emotion words differently. Havaldar et al. (2023b) analyzes Pride/Shame as a known cultural difference by prompting GPT-3.5 and GPT-4 to explore how these models handle pride and shame in the USA vs. Japan. They find that GPT-3.5 displays limited knowledge of culturally specific norms. Ahmad et al. (2024) expands the 19 cultural questions of the work Havaldar et al. (2023b) to 37 questions and evaluates ChatGPT for the low-resource Hausa language for sentiment analysis, but these evaluations are limited to 19/37 culturally relevant questions (not diverse and representative data), limited classes (Pride/Shame or positive/negative/mixed), and limited cultures (USA vs. Japan or Hausa).
@@ 3.1 Collecting Cultural Events
We manually craft scenarios, search on the web, and prompt LLMs to gather traditions, events, norms, and actions that elicit culturally different emotions across six target countries (UAE, USA, Germany, Ethiopia, India, and Mexico). We draw inspiration from the work of Havaldar et al. (2023b) to enhance topic diversity. Language representatives are asked to propose events distinct from those of other countries in the form of emotion-oriented questions. Importantly, these events do not contain explicit emotional keywords (as typically seen in traditional emotion datasets). We also refer to the International Survey on Emotion Antecedents and Reactions (ISEAR) (Scherer and Wallbott, 1994) data format, "When I … situations that cause a specific emotion"—a well-known English dataset for emotion analysis consisting of self-reported events from around 3,000 respondents across 37 countries and five continents. Table 1 displays the ten broad categories of the CuLEmo dataset and the number of questions in each category.
@@
Table 1: Ten broad categories and the number of events (# of Qn - number of questions) in the CuLEmo dataset.
@@ 3.2 Human-Adapted Translation
After collecting the events in English, we translate them into five target languages—Arabic, Amharic, German, Hindi, and Spanish—using Google Translate, followed by native-speaker approvals. Because the questions are simple “How do you feel when … ?” questions and lack explicit emotion keywords, translation quality did not affect their cultural content. While we acknowledge that existing works often depend on translating emotion-annotated data from English into other languages with their labels Kajava et al. (2020); Tahir et al. (2023); Bianchi et al. (2022), our translation process is done before any annotation. To ensure correctness, native speakers evaluate the translations. While most translations were acceptable, a few adjustments were made—e.g., to fix gender references and timing expressions for Amharic.
@@ 3.3 Language and Cultures Covered
Several factors guide our choice of languages, ensuring a broad range of cultural norms and concepts: (1) typological variety (five languages with four scripts), (2) geographical diversity (eastern vs. western contexts), (3) resource availability (low- vs. high-resource languages), and (4) the availability of native speakers for translation reviews.
@@ 3.4 CuLEmo Annotation
We use Amazon Mechanical Turk (MTurk) for most of our annotations, ensuring at least five native-speaker annotators per instance from the respective targeted country. We use a customized POTATO Pei et al. (2022) annotation tool for languages lacking sufficient MTurk annotators (e.g., Amharic) and recruit local native speakers who met our criteria. Annotators are fairly compensated \$12/hr. (Prolific [^3] https://www.prolific.com/’s minimum annotation wage is \$9/hr). In cases where no majority vote emerges among five annotators, we assign two additional annotations and use that majority vote.
Figure 2: Emotion label distribution across countries/languages: the number of instances in each emotion label across languages from a total of 400 events.
@@
Pairwise Label Agreements Across Countries We further examine label differences across countries by computing pairwise agreements after majority voting (Figure 3). Ethiopia and the United Arab Emirates exhibit the highest agreement at 55%, along with Germany and the United Arab Emirates; both exhibit a high neutral class (Figure 2), while Germany and India show the lowest agreement, at 29.0%. Labels from India thus diverge from those of other countries.
Figure 3: Pairwise emotion label agreements across countries/languages (ordered by their average agreement). Abbreviations: US = USA, MX = Mexico, DE = Germany, AE = UAE, ET = Ethiopia, and IN = India.
@@ 4.1 Task Formulation
To investigate the emotional comprehension capabilities of LLMs, we examine how they associate different cultures of a country with their corresponding languages. Specifically, we explore culture-aware emotion understanding via two main tasks: (1) emotion prediction and (2) sentiment analysis. All tested models are instruction-fine-tuned, except for the Aya-expanse model. We also experiment with prompts that do and do not include explicit country context, using the phrase "You live in <<country name>>," (where <<country name>> is one of the six targeted countries: UAE, USA, Ethiopia, Germany, India, and Mexico). Each task is framed as a text-generation problem, and the models are evaluated in a zero-shot setting.
@@ 4.2 Model Selection
We evaluate a variety of recent LLMs known for strong performance on standard benchmarks. We aim to include both smaller and medium-sized models, as well as open-source and proprietary models:
@@
    Open Source: LLaMA-3 (3.2-3B, 3.1-8B) Dubey et al. (2024); Meta AI (2024), Gemma (2B, 9B) Team Gemma et al. (2024), Aya (expanse-8b, 101-13B) Üstün et al. (2024); Dang et al. (2024), Ministral (3B, 8B) Mistral AI (2024)
@@
    Proprietary: GPT (3.5, 4) Achiam et al. (2024), Gemini-1.5 Team et al. (2024), Claude (3.5-sonnet, 3-opus) Anthropic AI (2024)
@@ 4.3 Multilingual Prompt Construction
To examine the impact of prompt language on model performance and to assess each model’s cultural awareness, we design both English and in-language prompts. The instruction, input text, and expected answer are in English when using English prompts. For in-language prompts, all elements are in one of the five target languages—Arabic (AR), Amharic (AM), German (DE), Hindi (HI), or Spanish (ES). Complete examples of our multilingual prompts for both emotion prediction and sentiment analysis are provided in Appendix C. We extract each model’s answer from its generated text using the PEDANTS tool Li et al. (2024).
@@ 5 Result and Analysis
Figure 4: Emotion prediction accuracy in radar chart across countries in English and in-language prompts. For lower-resource languages, English tends to work substantially better.
@@
Table 2: LLMs’ accuracy for the emotion prediction task. Columns labeled EN/AR/DE/AM/HI/ES show the prompt language for each corresponding culture (e.g., UAE is tested with English and Arabic). The highest-scoring model across English and in-language prompts is highlighted in bold, and the best model for the in-language prompt is underlined. * indicates a non-instruction fine-tuned model.
Questions Examples from CuLEmo US AE DE ET IN MX
@@
Table 3: Examples of emotion predictions from GPT-4 model. The first row under the country-code columns in each section represents the gold-label emotions, while the second row displays the predicted label. The predictions are using country context, by adding You live in <<country name>>, How do you feel when ... to the prompt.
English prompt In-language prompt
Lang. US AE DE ET IN MX AE DE ET IN MX
Llama-3.2-3B 0.44 0.41 0.36 0.35 0.31 0.46 0.13 0.19 0.30 0.25 0.37
Llama-3.1-8B 0.62 0.54 0.50 0.51 0.36 0.60 0.46 0.31 0.18 0.32 0.56
Gemma-2-2B 0.61 0.59 0.68 0.49 0.32 0.57 0.47 0.56 0.14 0.27 0.52
Gemma-2-9B 0.59 0.54 0.54 0.50 0.41 0.57 0.59 0.59 0.28 0.34 0.61
Aya-101-13B* 0.61 0.57 0.60 0.55 0.38 0.58 0.34 0.39 0.31 0.30 0.49
Ministral-8B 0.66 0.60 0.56 0.50 0.38 0.59 0.07 0.67 0.20 0.23 0.34
Claude-3.5-s. 0.56 0.51 0.39 0.49 0.36 0.60 0.51 0.35 0.53 0.38 0.61
Claude-3-op 0.52 0.40 0.37 0.49 0.36 0.59 0.45 0.30 0.47 0.37 0.58
Gemini1.5 0.53 0.51 0.39 0.48 0.39 0.59 0.54 0.42 0.52 0.37 0.65
GPT-4 0.58 0.55 0.46 0.48 0.37 0.63 0.46 0.45 0.27 0.41 0.64
Table 4: Emotion predicton results using English prompts and in-language prompts without specifying the country context, "You live in «country name»". The language column names are the two-letter targeted countries. The boldface result indicates the better results for each country. * indicates a non-instruction fine-tuned model.
5.1 Culture-Aware Emotion Prediction
Figure 5: Sentiment (positive, negative, and neutral) distribution across countries in the CuLEmo dataset.
LLMs USA UAE Germany Ethiopia India Mexico
EN EN AR EN DE EN AM EN HI EN ES
Llama-3.2-3B 0.57 0.41 0.15 0.20 0.24 0.48 0.51 0.60 0.56 0.55 0.60
Llama-3.1-8B 0.65 0.58 0.50 0.49 0.43 0.59 0.43 0.57 0.55 0.64 0.68
Gemma-2-2B 0.68 0.65 0.58 0.66 0.59 0.56 0.46 0.50 0.59 0.66 0.66
Gemma-2-9B 0.55 0.57 0.63 0.49 0.56 0.62 0.50 0.57 0.61 0.63 0.68
Aya-expanse-8b* 0.45 0.33 0.26 0.31 0.27 0.39 0.50 0.45 0.58 0.45 0.57
Aya-101-13B 0.65 0.64 0.55 0.69 0.45 0.55 0.53 0.52 0.56 0.61 0.63
Ministral-8B 0.70 0.66 0.06 0.60 0.73 0.59 0.46 0.54 0.24 0.65 0.53
Claude-3-opus 0.63 0.60 0.56 0.46 0.35 0.67 0.58 0.59 0.61 0.72 0.75
Gemini1.5-flash 0.63 0.63 0.63 0.48 0.50 0.63 0.60 0.60 0.53 0.69 0.71
GPT-4 0.67 0.61 0.54 0.48 0.51 0.62 0.49 0.58 0.62 0.72 0.75
Table 5: Accuracy for culture-aware sentiment analysis (positive/negative/neutral) with English and in-language prompts. The highest-scoring model across English and in-language prompts is highlighted in bold, and the best model for the in-language prompt is underlined. * is not an instruction fine-tuned model.
5.2 Culture Representation in LLMs
Results: According to Table 4, English prompt column category, the USA, Mexico, and Germany consistently achieve higher accuracy scores, while the UAE, Ethiopia, and India remain less accurately represented. This suggests that certain cultures may be more prevalent in the underlying training data.
5.3 Does Language Represent Country?
Can LLMs identify one country’s culture based solely on the prompt language? In this experiment, we remove explicit country context ("You live in <<country name>>") and test whether LLMs can infer cultural cues only from the language used in the prompt.
Results: Table 4, in-language prompt column category, shows that accuracy drops significantly when country context is omitted. Comparing these scores with Table 2 (where country context is included), we see consistent performance boosts (e.g. +1% in Spanish with GPT-4, +5% in Arabic with Gemini1.5, +6% in German with Ministral, +9% in Amharic with Claude-3.5-sonnet, and +21% in Hindi with GPT-4) when the prompt explicitly names the country. This indicates that language alone does not reliably convey cultural context. Notably, models like Claude-3.5-sonnet and GPT-4 show improvements in Indian and Ethiopian data when country context is specified, while English prompts (without “USA” context) are less affected. Overall, providing the country name remains crucial for accurate culture-aware emotion understanding, especially for less-resourced languages.
5.4 Culture-Aware Sentiment Analysis
For the sentiment analysis experiment, we follow De Bruyne et al. (2019) and Mostafazadeh Davani et al. (2022) by grouping emotions into positive (joy), negative (fear, anger, guilt, sadness), and neutral sentiments. Table 5 shows the accuracy of each LLM under this three-class setup.
Results: Table 5 shows better overall performance on sentiment analysis compared to fine-grained emotion classification. For example, GPT-4 gains notable accuracy (e.g., +22% for Hindi). Similarly, the highest score appears for Mexican culture (Spanish) with Claude-3-opus and GPT-4, each at 75%. By contrast, Aya-expanse-8b struggles more, as it is not instruction fine tuned. Smaller models like Gemma-2-2B are competitive with proprietary models. Still, performance drops persist for Amharic and Hindi, reflecting the challenges of culture-aware tasks in lower-resource contexts.